OpenStudy (anonymous):

What percent of 67 is 8

jhonyy9 (jhonyy9):

67 ----- 100% 8 ---- x --------------- x = 8*100 /67 = ? hope you can calculi it ,yes ?

jhonyy9 (jhonyy9):

x=800/67 = 11,94 %